the locus of points from which length of tangent to the two circles x²+y²+4x+3=0 x²+y²-6x+5=0 are in the ratio 2:3 is a circle with centre

Grade:12

1 Answers

Manu Singh
6 years ago
The length of a tangent from any point (x1,y1) to a circle S=0 is √S1.Let the point be (x,y).Thus According to Question√x^2+y^2+4x+3/√x^2+y^2-6x+5 =2/3On solving further we get,x^2+y^2+12x+7/5=0The center of this circle is (-6,0)